# Low energy excitations in A-site ordered SmBaMn2O6

### v1

@@ Abstract
The electron in a solid can be considered a bound state of the three independent, fundamental degrees of freedom creating quasi-particles: spinons, carrying the electron spin; plasmons carrying the collective charge mode and orbitons carrying its orbital degree of freedom. These fundamental degrees of freedom could form ordering states in which dynamics or collective motions could occurr and manifest as low-energy excitations. The exotic properties that appear in the materials exhibiting these electronic orderings are associated with these low-energy excitations. Although the orbital order (OO) and its coupling to the spin system creates very interesting phenomena, the microscopic origin of OO has been much less explored than other electronic properties as it is very difficult to directly access experimental information from OO. Due to the recent improvement in energy resolution and flux, soft x-ray resonant inelastic scattering (RIXS) allows for a re-examination of orbital excitations in manganites. Here, we present a study of low energy excitations in half doped A-site ordered SmBaMn2O6 through a combination of RIXS and soft x-ray resonant elastic scattering (REXS) measurements. The obtained experimental data confirm the OO at = (0.25, 0.25, 0) and find various low energy excitations below 200 meV. while several excitations can be assigned to be of magnetic and phononic origin, a group of excitations between 80 and 200 meV show a temperature dependence distinctively following that of the OO making them possible candidates for orbital excitations.
@@ I Introduction
In the last years, resonant elastic x-ray scattering (REXS) has provided more information about orbital physics by directly observing the orbital order in different manganite compounds[5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. However, the dynamics of the orbital degree of freedom is still far from understood. Several theoretical studies have predicted that orbital excitations are distinguishable by characteristic variations in the RIXS amplitude as a function of incident energy and momentum transfer[15, 16, 17] from other excitations in the solid. These studies predict that both single- and double-orbiton excitations are allowed, with intensities that are of the same order. In one-dimensional cuprates, orbitons are one of the three quasiparticles that electrons in solids are able to split into during the process of spin-charge separation. In two-dimensional cuprates, pure dispersive orbital excitations have been observed by RIXS [18]. However, the detection and study of the dispersion of orbitons in one of the prototypical orbital ordered materials, i.e. half doped manganites has been elusive so far. To our knowledge only K-edge RIXS studies of doped manganites can be found in the literature to date[19, 20, 21, 22]. As RIXS experiments at the K-edge probe the 3d states indirectly[22], studies at the Mn -edge of doped manganites with high energy resolution are required to bring light into this issue.
@@
The half doped SmBaMn2O6 (SBMO) manganite crystallizes in the slightly distorted perovskite structure ABO3 depicted in Figure 1(a). In this compound the MnO2 octahedra are sandwiched by two distinct layers built of SmO and BaO[23] in contrast to the disordered Sm0.5Ba0.5MnO3 (Figure 1(b)) with a random distribution of Sm and Ba ions. The A-site ordered RBaMn2O6 compounds with R=Sm, Eu, Gd, Tb, Dy, Ho and Y undergo a two-step charge-orbital order (COO) transition at temperatures Tco1 = 380 K and Tco2 = 210 K. Figure 1c shows magnetization measurements of SBMO single crystal where these transitions can be identified. Below Tco2 = 210 K, a disappearance of the simple perovskite four-fold periodicity along the c-axis has been suggested to be caused by the re-stacking of the COO planes. The exact model, however, remains controversial[24, 25, 26, 27, 28]. The issue of the unresolved models arises partly from the fact that most experimental studies were performed on polycrystalline samples which limits the direct observation of its microscopic physical properties. A recent convergent beam electron diffraction experiment on a single crystal of A-site ordered SBMO determined that the space group for the Tco2-COO and Tco1-COO phases are P21am and Pnam respectively [29]. This space group change suggests that an even more complex structural arrangement occurs below Tco2 as a consequence of the stacking of the COO. Some theoretical studies even predict a ferroelectric state in analogy to Pr(Sr0.1Ca0.9)2Mn2O7 [29, 30]. The existence of this second charge order transition at 210 K, unique for the A-site ordered manganites is responsible for a distinct temperature dependence of the OO parameter that will be discussed later.
@@ II Experimental details
The RIXS and REXS experiments were conducted at the I21-RIXS beamline at Diamond Light Source, United Kingdom[31]. The sample was mounted with the [110] plane lying in the scattering plane onto a copper sample holder fixed to a liquid He flow cryostat which achieves temperatures between 7 K and 380 K. Experiments were performed using linear horizontal or vertical polarized light leading to or incident photon polarization in respect to the horizontal scattering geometry.
We tuned the incident photon energy to the resonance of the Mn L2,3 absorption edges for the RIXS measurements. The total energy resolution achieved was about 24 meV FWHM. RIXS and REXS signals were collected without polarization analysis. For all RIXS spectra, the elastic (zero-energy loss) peak positions were determined by the elastic scattering spectrum from carbon tape placed near the sample surface and then fine adjusted by its Gaussian fitted elastic peak position. The Miller indices in this study are defined with reference to the crystal structure x x 2 with being the cubic perovskite unit cell. The experimental geometry of the sample is depicted in Figure 1(d). The temperature of the measurements was 15 K unless otherwise stated. The REXS measurements were performed using a rotating photodiode inside the sample vessel. The RIXS measurements were collected through continuous rotation of the spectrometer arm due to the 3D nature of the investigated compound. The diffraction peaks have been analyzed using Pseudo-Voigt functions while the inelastic peak were described using Gaussian functions.
@@ III Results and discussion
Figure 2 shows the REXS data of energy dependence of the orbital (1/4 1/4 0) reflection in the vicinity of the Mn edges. The insert of Figure 2 shows the comparison to the experimental x-ray absorption (XAS). Clearly, the resonant reflection intensity across the Mn edges is distinct from the XAS signal and evidences the OO [10, 11, 14]. More specifically, the present measurements on SBMO shows the same differences with the layered La0.5Sr1.5MnO4 spectra observed previously in literature [10]. These differences confirm that the OO of the eg electrons is of x2-z2/y2-z2 type in A-site ordered SBMO, while a 3x2-r2/3y2-r2 type of orbital order is present in the layered manganite.
@@
Following the REXS measurements, we performed RIXS measurements for various incident energies with the and 2 angles ensuring a fixed momentum transfer at the (1/4 1/4 0) reflection when varying the x-ray energy. The integrated intensities between 0 and 100 meV of these RIXS spectra are shown as red dots in Figure 2 together with the REXS signal. The measurements are plotted in different y-axis for clarity. Both RIXS and REXS signals are very similar confirming the excellent alignment of the spectrometer arm to = (0.25, 0.25, 0).
@@
Figure 3 shows interesting information regarding the collected RIXS spectra, the difference of spectral weight between the measurements done at the and edges. Due to the self absorption, while the RIXS spectra shows more clearly the d-d and low energy excitations, the spectra is mainly dominated by the elastic signal. Keeping this information in mind we proceeded to study the low-energy excitations in order to explore the potential existence of the collective orbiton.
@@
We now discuss the inelastic excitation at the OO OO =(0.25, 0.25, 0). Given that the inelastic signal is stronger at the Mn edge compared to the edge, we collected the momentum-dependent RIXS spectra across the OO scattering peak at the Mn edge. Figures 4a and 4b, show the momentum-dependent RIXS maps probed by the linear and incident polarizations, respectively. However, the OO elastic scattering signal completely dominates the spectra with no appreciable spectral weight from the low energy excitations. The low energy excitation spectra between 0 and 200 meV at three different positions (0.25, 0.25, 0), (0.245, 0.245, 0) and (0.255, 0.255, 0) and T=15K are shown in Figure 5. In the normalized plot (upper insert) weak low energy excitations can be observed at around 80 meV when is not exactly at the OO Bragg condition. To examine further whether there are any detectable low-energy excitations resonating with the OO peak, we show in the lower inset of Figure 5 the normalized RIXS spectra with incident and polarizations together with the elastic scattering spectrum from a carbon reference sample collected in the same geometry. The nearly identical spectra between the manganite and the reference sample suggest that any possible low-energy excitations are shadowed by the OO scattering at this wave vector.
The RIXS spectra were collected both at the OO condition and off this condition in order to extract the intensity of the low energy excitations. The offsets in tilt and angles were increased to achieve an alignment of . RIXS spectra were collected for the same energy values as the data presented in Figure 2, while varying the experimental geometry to ensure is preserved for each energy value. The integrated intensity of these RIXS spectra is shown in the insert of figure 2 together with the XAS signal.
@@
In addition to the magnons, we expect also phonons to be present in this range of energy. Akahoshi et al. [28] studied SBMO using optical conductivity and Raman scattering. These techniques are restricted to collect data at = 0 and are not sensitive to the dispersion of the observed excitations. Two major peaks are observed in the Raman spectra around 500 and 620 cm-1 (62 and 77 meV) that are assigned to the Jahn-Teller (JT) and breathing modes respectively. These phonon energies are displayed in Figure 6 as dashed grey lines.
@@
Several works have considered orbital excitations in the parent compound LaMnO3[34, 35, 36, 17], with significant variation of the role attributed to phonons. Van der Brink[34] used self-consistent second order perturbation theory in the orbiton-phonon coupling and found that the orbiton dispersion is strongly reduced by the electron-phonon coupling. In addition, he found that this coupling also mixes the orbiton and phonon modes and causes satellite structures in the orbiton and the phonon spectral function. Schmidt et al. [36] performed a calculation to provide information on the nature of orbital excitations in the presence of substantial orbiton phonon coupling. In their model, they considered quantitatively the dynamic orbiton-phonon interaction that corresponds to the creation or annihilation of a local distortive phonon in the presence of an orbiton. Their results show that the system becomes more local when increasing the orbiton-phonon coupling, i.e., the effective bandwidth of the orbiton band decreases as the orbiton-phonon coupling increases. In a later work, that focussed completely on the observation of orbital excitations in LaMnO3 using RIXS, Forte et al.[17] used the ultra short core-hole lifetime expansion for RIXS and obtained a gapped orbiton spectrum. Since their calculated orbital Hamiltonian did not have a continuous symmetry, the Goldstone models were absent. In addition they found that at high symmetry points in the Brillouin zone, the intensity of specific orbiton branches vanishes. This result is consistent with our experimental data that shows that at the OO vector the OO diffraction peak dominates the spectra. Unfortunately, we are not able to fully separate the contribution of magnons, phonons, and orbitons to the observed spectra. Further experiments using polarization analysis would be required to obtain a clearer picture of the nature of a possible orbiton dispersion in half doped manganite SBMO.
We are aware that these experimental observations are not sufficient to univocally claim that the observed excitations correspond to the elusive orbitons. However, the observed correlation is a strong motivation for further studies. We believe that in order to properly explore the low energy excitations in half doped manganites, further improved energy resolution and additional polarization analysis of the scattered radiation would be essential to clarify the electronic nature of the observed excitations.

### v2

# apsrev42Control Low energy excitations in A-site ordered SmBaMn2O6

@@ Abstract
The electron in a solid can be considered a bound state of the three independent, fundamental degrees of freedom creating quasi-particles: spinons, carrying the electron spin; plasmons carrying the collective charge mode and orbitons carrying its orbital degree of freedom. These fundamental degrees of freedom could form ordering states in which dynamics or collective motions could occur and manifest as low-energy excitations. The exotic properties that appear in the materials exhibiting these electronic orderings are associated with these low-energy excitations. Although the orbital order (OO) and its coupling to the spin system creates very interesting phenomena, the microscopic origin of OO has been much less explored than other electronic properties as it is very difficult to directly access experimentally. Owing to the recent improvement in energy resolution and flux, soft x-ray resonant inelastic scattering (RIXS) allows for a re-examination of orbital excitations in manganites. Here, we present a study of low energy excitations in half doped A-site ordered SmBaMn2O6 through a combination of RIXS and soft x-ray resonant elastic scattering (REXS) measurements. We confirm the existence of OO at = (0.25, 0.25, 0) and find various low energy excitations below 200 meV. While several excitations can be assigned to be of magnetic and phononic origin, a group of excitations between 80 and 200 meV show a temperature dependence closely following that of the OO making them possible candidates for orbitons.
@@ I Introduction
In the last years, resonant elastic x-ray scattering (REXS) has provided much information about orbital physics by directly observing the orbital order in different manganite compounds[23, 36, 31, 9, 30, 12, 14, 8, 29, 13]. Several theoretical studies have predicted that orbital excitations are distinguishable by characteristic variations in the RIXS amplitude as a function of incident energy and momentum transfer[19, 18, 11] from other excitations in the solid. These studies predict that both single- and double-orbiton excitations are allowed, with intensities that are of the same order. In one-dimensional cuprates, orbitons are one of the three quasiparticles that electrons in solids are able to split into during the process of spin-charge separation. In two-dimensional cuprates, pure dispersive orbital excitations have been observed by RIXS [27]. However, the detection and study of the dispersion of orbitons in one of the prototypical orbital ordered materials, i.e. half doped manganites has been elusive so far. To our knowledge only K-edge RIXS studies of doped manganites can be found in the literature to date[22, 17, 21, 15]. As RIXS experiments at the K-edge probe the 3d states indirectly[15], studies at the Mn -edge of doped manganites with high energy resolution are required to bring light into this issue.
@@
The half doped SmBaMn2O6 (SBMO) manganite crystallizes in the slightly distorted perovskite structure ABO3 depicted in the insert of Figure 1(a). In this compound the MnO2 octahedra are sandwiched by two distinct layers built of SmO and BaO[4] in contrast to the disordered Sm0.5Ba0.5MnO3 with a random distribution of Sm and Ba ions. The A-site ordered RBaMn2O6 compounds with R=Sm, Eu, Gd, Tb, Dy, Ho and Y undergo a two-step charge-orbital order (COO) transition at temperatures Tco1 = 380 K and Tco2 = 210 K hile this transition is absent for the disordered compounds. Figure 1(a) shows magnetization measurements of SBMO single crystal where these transitions can be identified. Below Tco2 = 210 K, a disappearance of the simple perovskite four-fold periodicity along the c-axis has been suggested to be caused by the re-stacking of the COO planes. The exact model, however, remains controversial[3, 2, 24, 7, 5]. The issue of the unresolved models arises partly from the fact that most experimental studies were performed on polycrystalline samples which limits the direct observation of its microscopic physical properties. A recent convergent beam electron diffraction experiment on a single crystal of A-site ordered SBMO determined that the space group for the Tco2-COO and Tco1-COO phases are P21am and Pnam respectively [1]. This space group change suggests that an even more complex structural arrangement occurs below Tco2 as a consequence of the stacking of the COO. Some theoretical studies even predict a ferroelectric state in analogy to Pr(Sr0.1Ca0.9)2Mn2O7 [1, 32]. The existence of this second charge order transition at 210 K, unique for the A-site ordered manganites is responsible for a distinct temperature dependence of the OO parameter that will be discussed later.
@@ II Experimental details
The RIXS and REXS experiments were conducted at the I21-RIXS beamline at Diamond Light Source, United Kingdom[37]. The sample was mounted with the [110] plane lying in the scattering plane onto a copper sample holder fixed to a liquid He flow cryostat which achieves temperatures between 7 K and 380 K. Experiments were performed using linear horizontal or vertical polarized light leading to or incident photon polarization in respect to the horizontal scattering geometry. The REXS measurements were performed using a photodiode as detector without the possibility to differentiate the energy of the outgoing photons.
We tuned the incident photon energy to the resonance of the Mn L2,3 absorption edges for the RIXS measurements. The total energy resolution achieved was about 24 meV FWHM. RIXS and REXS signals were collected without outgoing x-ray polarization analysis. For all RIXS spectra, the elastic (zero-energy loss) peak positions were determined by the elastic scattering spectrum from carbon tape placed near the sample surface and then fine adjusted by its Gaussian fitted elastic peak position. The Miller indices in this study are defined with reference to the crystal structure x x 2 with being the cubic perovskite unit cell. The experimental geometry of the sample is depicted in Figure 1(d). The temperature of the measurements was 15 K unless otherwise stated. The REXS measurements were performed using a rotating photodiode inside the sample vessel. The RIXS measurements were collected through continuous rotation of the spectrometer arm due to the 3D nature of the investigated compound. The diffraction peaks have been analyzed using Pseudo-Voigt functions while the inelastic peak were described using Gaussian functions.
@@ III Results and discussion
Figure 2 shows the REXS data of energy dependence of the orbital (1/4 1/4 0) reflection in the vicinity of the Mn edges. The insert of Figure 2 shows the comparison to the experimental x-ray absorption (XAS). Clearly, the resonant reflection intensity across the Mn edges is distinct from the XAS signal and evidences the OO [12, 14, 13]. More specifically, the present measurements on SBMO shows the same differences with the layered La0.5Sr1.5MnO4 spectra observed previously in literature [12]. These data confirm that the OO of the eg electrons is of x2-z2/y2-z2 type in A-site ordered SBMO, while a 3x2-r2/3y2-r2 type of orbital order is present in the layered manganite.
@@
Following the REXS measurements, we performed the energy dependence RIXS measurements with the and 2 angles ensuring a fixed momentum transfer at the (1/4 1/4 0) reflection when varying the x-ray energy. The integrated intensities between 80 meV of these RIXS spectra are shown as red dots in Figure 2 together with the REXS signal. Both RIXS and REXS signals are very similar confirming the excellent alignment of the spectrometer to = (0.25, 0.25, 0). In order to minimize the elastic signal, the RIXS measurements shown later were collected both at and off from the OO Bragg condition at . In the insert of Figure 2 the comparison with the observed energy dependence measured at fixed with the REXS and the XAS signal is shown.
@@
Figure 3 shows the difference of spectral weight between the measurements done at the and edges. Due to the self absorption, the RIXS spectra shows more clearly the d-d and low energy excitations, whereas the spectra is mainly dominated by the elastic signal. Keeping this information in mind we proceeded to study the low-energy excitations at the incident energy of 652.85 eV in order to explore the potential existence of the collective orbiton.
@@
We now discuss the inelastic excitation at the OO OO =(0.25, 0.25, 0). Given that the inelastic signal is stronger at the Mn edge compared to the edge, we collected the momentum-dependent RIXS spectra across the OO scattering peak at the Mn edge. Figures 4a and 4b, show the momentum-dependent RIXS maps probed by the linear and incident polarizations, respectively. However, the OO elastic scattering signal completely dominates the spectra with no appreciable spectral weight from the low energy excitations and no dependence with the incident polarization is observed. This absence of polarization dependence of the OO Bragg peak is consistent with the previous REXS studies of OO in powder SBMOGarcía-Fernández et al. [12, 14, 13].
Figure 5 shows the normalized low energy low energy excitation spectra between 0 and 200 meV at three different positions (0.25, 0.25, 0), (0.245, 0.245, 0) and (0.255, 0.255, 0) collected at T=15K and with linear horizontal incident polarization These positions correspond to the grey dashed lines in 4a and are shown together with the normalized elastic scattering spectrum from a carbon reference sample measured at the same geometry to illustrate the achieved energy resolution. The imperfect tail off the carbon tape gaussian distribution reflects that the instrumental resolution function is distorted by the optics aberration. The near perfect gaussian shape of the OO condition is due to the strong Bragg diffraction which suppresses the aberration. When is not exactly at the OO Bragg condition, weak low energy excitations can be observed at around 80 meV. These data suggest that any possible low-energy excitations are shadowed by the OO scattering at this wave vector.
(b) Incident polarization.
@@
In addition to the magnons, we expect also phonons to be present in this range of energy. Akahoshi et al. [5] studied SBMO using optical conductivity and Raman scattering. These techniques are restricted to = 0 and are not sensitive to the dispersion of the observed excitations. Two major peaks are observed in the Raman spectra around 500 and 620 cm-1 (62 and 77 meV) that are assigned to the Jahn-Teller (JT) and breathing modes respectively. These phonon energies are displayed in Figure 6 as dashed grey lines.
@@
Several works have considered orbital excitations in the parent compound LaMnO3[35, 6, 28, 11], with significant variation of the role attributed to phonons. Van der Brink[35] used self-consistent second order perturbation theory in the orbiton-phonon coupling and found that the orbiton dispersion is strongly reduced by the electron-phonon coupling. In addition, he found that this coupling also mixes the orbiton and phonon modes and causes satellite structures in the orbiton and the phonon spectral function. In this context the observed excitations between 80 and 200 meV could be explained as the orbital collective excitations plus these phonon satellites whose energy would involve contributions from both the orbital and the lattice components.
Schmidt et al. [28] performed a calculation to provide information on the nature of orbital excitations in the presence of substantial orbiton phonon coupling. In their model, they considered quantitatively the dynamic orbiton-phonon interaction that corresponds to the creation or annihilation of a local distortive phonon in the presence of an orbiton. Their results show that the system becomes more local when increasing the orbiton-phonon coupling, i.e., the effective bandwidth of the orbiton band decreases as the orbiton-phonon coupling increases. In a later work, that focussed completely on the observation of orbital excitations in LaMnO3 using RIXS, Forte et al.[11] used the ultra short core-hole lifetime expansion for RIXS and obtained a gapped orbiton spectrum. Since their calculated orbital Hamiltonian did not have a continuous symmetry, the Goldstone models were absent. In addition they found that at high symmetry points in the Brillouin zone, the intensity of specific orbiton branches vanishes. This result is consistent with our experimental data that shows that at the OO vector the OO diffraction peak dominates the spectra.
We are aware that these experimental observations are not sufficient to univocally claim that the observed excitations correspond to the elusive orbitons. However, the observed correlation is a strong motivation for further studies. We believe that in order to properly explore the low energy excitations in half doped manganites, further improved energy resolution and additional polarization analysis of the scattered radiation would be essential to clarify the electronic nature of the observed excitations. As it would allow us to distinguish the excitations that are present in each of the four polarization channels. Note that for coupled orbital-phonon modes, one can only see the ’orbital’ contributions as is the case for the measurement of phonons at the oxygen K-edge, see e.g. Ueda et al.Ueda et al. [33].